OpenStudy (anonymous):

Marie's bakery has 8 different cupcakes to put in the display case. If the cupcakes are randomly placed in the display case to form a single row, how many different ways can the 8 cupcakes be put in the display case? A. 6! B. 7! C. 8! D. 9!

OpenStudy (anonymous):

!<~~~ what does this mean???

OpenStudy (anonymous):

its a factorial. 1! = 1 2! = 1*2 3! = 1*2*3 4! = 1*2*3*4 and so on... the answer is C) 8!